OpenStudy (fairegaming):

**Answering 6-9th Grade Math** *POST BELOW* Hey! I may not be allowed to do this, but eh worth a try, I am answering 6-9th grade questions easily, "No Medal Required" unless I really helped ;)

OpenStudy (anonymous):

Can you help me with this question? A. -9 ÷5/8= 1. - 8/45 B. 5/75 C. -14 2/5 D. 5 5/8 I don't understnad it :<

OpenStudy (fairegaming):

A then 1 then B, C, D? Whaaaaa. Wheres the Question.

OpenStudy (sloppycanada):

I assume it's matching the 1 to A.

OpenStudy (fairegaming):

Maybe, or they want you to answer A, 1, B, C, D but I doubt it.

OpenStudy (sloppycanada):

No it's a matching thing.

OpenStudy (anonymous):

Sorry I did not realize I did that V V sorry. The question is: -9 ÷5/8= One of them are A. - 8/45 B. 5/75 C. -14 2/5 D. 5 5/8

OpenStudy (fairegaming):

Then wheres B, 2?

OpenStudy (fairegaming):

Yeah I thought something was... Fishy :P

OpenStudy (fairegaming):

You need "Help" or the "Answer"?

OpenStudy (anonymous):

Either is fine

OpenStudy (sloppycanada):

I'm going to have to disagree with you there.

OpenStudy (anonymous):

um ok then

OpenStudy (anonymous):

Sorry...

OpenStudy (sloppycanada):

Gonna have to go with C.

OpenStudy (fairegaming):

Hmm.. How is that so?

OpenStudy (sloppycanada):